# Towards Optimal Distributed Edge Coloring with Small Palettes

### v1

@@ 4.3 Phase 2: Assign two exclusive edges to each cluster
Next, we rearrange the edges of $M$ in order to assign two exclusive edges to each cluster while maintaining the matching property. The two exclusive edges are needed in the subsequent steps to adjust the coloring of the intercluster edges such that the coloring can be completed within the cluster. This step is based on the HSO subroutine stated in Theorem B.8 and we describe in the following the specific instance that we use in our algorithm.
Algorithm 3 Flip$(\mathcal{C},M)$
1:Input: a $\Delta$-regular graph $G=(V,E)$, a clustering $\mathcal{C}$ and a matching $M\subseteq E$.
2:Output: a matching $M^{\prime}\subseteq E$ and an assignment $\psi:M^{\prime}\to\mathcal{C}^{\prime}$
3:
4:Each cluster $C^{-1}(r)\in\mathcal{C}^{\prime}$ finds an integer $k>0$ and a set of vertices $S\subseteq V(C^{-1}(r))$ of size at least $4\cdot\Delta^{3}$ such that $\mathrm{dist}(r,v)=k$ for all $v\in S$. $\triangleright$ Lemma 4.7
5:Construct auxiliary hypergraph $H=(\mathcal{C}^{\prime},M)$ with $\delta_{H}>\Delta\cdot r_{H}$. $\triangleright$ Lemma 4.10
6:Compute a $2$-HSO $\psi$ in $H$triangleright Lemma 4.12
7:Move matching edges closer to their cluster $\triangleright$ Lemma 4.14
8:return $M^{\prime},\psi$
The high level idea is that each cluster proposes to nearby edges in the existing maximal matching $M$. Each matching edge accepts exactly one proposal and we swap the edge towards the cluster that won it. This process can be modeled as an HSO instance where each matching edge in $M$ represents an hyperedge that contains as vertices all clusters that sent a proposal to that edge. Since a directed hyperedge $e$ is outgoing for exactly one of its vertices $v\in e$, we will say that $v$ won the proposal for $e$. However, the hypergraph sinkless orientation problem only guarantees at least one outgoing hyperedge for each vertex. Therefore, in order to achieve two exclusive hyperedges for each cluster, we split each cluster into two virtual nodes and let both participate in the HSO independently. In the following we formalize this idea and prove that this provides the desired properties as well as a deterministic runtime of $\mathcal{O}(\log n)$ rounds. First, we let each cluster $C^{-1}(r)\in\mathcal{C}$ that is not already edge-degree choosable select a subset of vertices $S\subseteq V(C^{-1}(r))$ to send proposals to a matching edges in their neighborhood. Let $\mathcal{C}^{\prime}$ denote the set of clusters, which are not already edge-degree choosable. Crucially, in order to satisfy the colorful condition from Lemma 3.4 we require that the vertices in $S$ are all leaves in the spanning tree $T(r)\subseteq C^{-1}(r)$ at the same distance from the root $r$.
@@
For each cluster $C^{-1}(r)\in\mathcal{C}^{\prime}$, there exists a $k>0$ and a set of vertices $S\subseteq V(C^{-1}(r))$ of size at least $4\cdot\Delta^{3}$ such that
  • $\mathrm{dist}(r,v)=k$ for all $v\in S$ and
  • $v$ is a leaf in $T(r)$ for all $v\in S$.
@@
To compute an orientation with two outgoing edges per cluster we simply split each vertex into two virtual nodes with half the degree each. Thus, the minimum degree of the resulting hypergraph $H^{\prime}$ is still at least $\Delta^{2}$. Therefore, we can solve HSO in $\mathcal{O}(\log_{\delta/r}n)=\mathcal{O}(\log_{\Delta}n)$ rounds deterministically and in $\mathcal{O}(\log_{\delta/r}\delta+\log_{\delta/r}\log n)=\mathcal{O}(\log_{\Delta}\log n)$ rounds randomized using the algorithms from Theorem B.8 and Theorem B.9.
@@ 4.4 Phase 3: Switch colors in order to complete the coloring
Now that every cluster got access to two exclusive edges, which are at the same distance from the central vertex of the cluster, we show that we can modify these colors in order to satisfy the colorful condition from Lemma 3.4 for each cluster. To show that we can perform this procedure for all clusters in parallel, we argue that the colorful condition can still be fulfilled, even if the colors of all edges not adjacent to the two exclusive edges are modified adversarially.
@@
Recall that $V_{k}:=\{v\in V_{T}:\mathrm{dist}_{T}(v,r)=k\}$. Since $u$ and $u^{\prime}$ are leaves in $T$ they are incident to exactly $\Delta-1$ edges in $E\setminus E_{T}$. Therefore, $\varphi(N_{E}(V_{k}))$ already contains at least $\Delta-1$ distinct colors. If $\lvert\varphi(N_{E}(V_{k}))\rvert\geq\Delta$ we simply set $\varphi^{\prime}=\varphi$. Otherwise, if $\lvert\varphi(N_{E}(V_{k}))\rvert=\Delta-1$, we set $\varphi^{\prime}(uv)=2\Delta-2$. Since $\varphi(e)<2\Delta-2$ for all $e\in E_{k}$ we only need to argue that the color $\varphi(uv)$ still remains in $\varphi^{\prime}(E_{k})$. Consider the sets $N:=\{uw:uw\in E\setminus E_{T}\}$ and $N^{\prime}:=\{u^{\prime}w:u^{\prime}w\in E\setminus E_{T}\}$. Since $u,u^{\prime}$ are both leaves in $T$, both $N$ and $N^{\prime}$ have cardinality $\Delta-1$. Moreover, since $\lvert\varphi(E_{k})\rvert=\Delta-1$ and any two edges in $N$ (or $N^{\prime}$) are adjacent, we have that $\varphi(N)=\varphi(N^{\prime})$. Hence, there exists an edge $u^{\prime}v^{\prime\prime}$ such that $\varphi(uv)=\varphi(u^{\prime}v^{\prime\prime})=\varphi^{\prime}(u^{\prime}v^{\prime\prime})\in\varphi^{\prime}(E_{k})$. Thus, we have $\varphi^{\prime}(E_{k})=\Delta$. Note that this argument holds even if the colors of the edges not adjacent to $u$ or $v$ are modified adversarially.
Proof 4.18 (Proof of Theorem 1.1).
@@
For $\Delta=\mathcal{O}(1)$ we can simply apply the $(3\Delta/2)$-edge coloring algorithm from Theorem B.12 to get a $(2\Delta-2)$-edge coloring in $\mathcal{O}(\log n)$ rounds. Hence, for the remainder of the proof we can assume $\Delta$ to be larger than a sufficiently large constant. Computing an MIS on the power graph $G^{8}$ takes $\widetilde{\mathcal{O}}(\log^{5/3}n)$ rounds using the algorithm from Theorem B.10. Next, the clustering procedure can be done in a constant number of rounds, since the radius of each cluster is at most 8. For the $(2\Delta-3)$-edge coloring $\varphi$ of $G-E_{T}$, as well as the maximal matching, we can again use the algorithm from Theorem B.10. Then, according to Section 4.3 we can rearrange the matching edges in a way that each cluster receives two exclusive edges to recolor. The rearrangement can be modeled as an instance of HSO with minimum degree $\frac{\Delta^{3}}{4}$ and maximum rank $\Delta^{2}$. The vertices of the virtual hypergraph are the clusters in $G$. Since each cluster has constant diameter, we can simulate the hypergraph with just a constant overhead factor. Thus, the deterministic HSO algorithm from Theorem B.8 runs in $\mathcal{O}(\log_{\Delta}n)$ rounds. Next we apply Lemma 4.16 to ensure that for each cluster there exists a $k$ such that $\lvert\varphi(N_{E}(V_{k}))\rvert\geq\Delta$ and by Lemma 4.16 we can extend $\varphi$ to a proper $(2\Delta-2)$-edge coloring of $G$.
We note that since the current state-of-the-art complexity for $(2\Delta-1)$-edge coloring and MIS coincide, this simple version of our algorithm already yields the $\widetilde{\mathcal{O}}(\log^{5/3}\log n)$ round algorithm claimed in Corollary 1.2. Our optimized version of the algorithm actually does not improve on the total runtime, but replaces the MIS and maximal matching subroutines by ruling sets and edge-ruling sets, respectively. Since $(2\Delta-1)$-edge coloring is in general a much easier problem than MIS, we expect that this optimization might lead to better runtimes in the future.
4.5 Removing MIS from our reduction
@@
As it turns out, computing an MIS is not strictly necessary for our reduction. Indeed, using a $(2,k)$-ruling set just increases the diameter of our clusters. By tuning the parameter $k$ one finds that a $(2,\mathcal{O}(\log\Delta))$-ruling set is optimal for our purposes. This change only affects the runtime of some of our subroutines. For the second change, where we substitute the maximal matching by a $2$-edge ruling set, we actually need a slightly generalized version of how we set up our auxiliary hypergraph. We defer this modification to Section A.1 in the appendix. The changes to the simple version of Section 4.1 are also highlighted red in Algorithm 4.
@@
Algorithm 4 Optimized $(2\Delta-2)$-edge coloring (high level overview)
1:Compute a $(2,\mathcal{O}(\log\Delta))$-ruling set $\mathcal{R}$ on $G^{8}$triangleright Corollary B.3
@@
4:Compute a $2$-edge ruling set $E_{\mathcal{R}}$ of $T$triangleright Theorem B.7
5:Compute $2$-HSO $\psi:\mathcal{C}\to E_{\mathcal{R}}$ in $H=(\mathcal{C},E_{\mathcal{R}})$triangleright Theorem B.8
@@
8:     Obtain $\varphi^{\prime}$ from $\varphi$ by changing the colors of the edges in $\chi(\psi(\mathfrak{C}))$triangleright Lemma 4.16
@@
10:Extend $\varphi^{\prime}$ to a $(2\Delta-2)$-edge coloring $\varphi^{\star}$ of $G$triangleright Lemma 3.4
@@
Proof 4.19.
The $(2,\mathcal{O}(\log\Delta))$-ruling set can be computed in $\mathcal{O}(\log\Delta+\log^{\ast}n)$ rounds according to Corollary B.3. As an immediate consequence, the diameter of our clusters now increases to $\mathcal{O}(\log\Delta)$. This implies that simulating the hypergraph now incurs a $\mathcal{O}(\log\Delta)$ factor overhead. Therefore the complexity of computing a HSO increases to $\mathcal{O}(\log_{\Delta}n\cdot\log\Delta)=\mathcal{O}(\log n)$ rounds. The $(2\Delta-3)$-edge coloring of $G-E_{T}$ remains unchanged. Further, we can also relax the requirements for our matching $M$ to $2$-edge ruling set, which can be computed in $\mathcal{O}(\log^{\ast}n)$ rounds. This requires us to be a little more careful in how we set up our hypergraph, which we defer to Section A.1 in the appendix. Therefore the total runtime of our reduction works out to be
@@ 5 Randomized $(2\Delta-2)$-edge coloring algorithm
In order to achieve a $\mathcal{O}(\log\log n)$-round randomized reduction, we only need to replace the ruling set algorithm by a faster randomized variant in Corollary B.5 and the deterministic HSO algorithm by its randomized counterpart in Theorem B.9. The changes to the optimized deterministic version are also highlighted red in Algorithm 5
@@
1:Compute a $(2,\mathcal{O}(\log\Delta)\cap\mathcal{O}(\log\log n))$-ruling set $\mathcal{R}$ on $G^{8}$triangleright Corollary B.5
@@
3:Compute a $(2\Delta-3)$-edge coloring $\varphi$ of $G-E_{T}$triangleright Corollary B.11
4:Compute a $2$-edge ruling set $E_{\mathcal{R}}$ of $T$triangleright Theorem B.7
5:Compute $2$-HSO $\psi:\mathcal{C}\to E_{\mathcal{R}}$ in $H=(\mathcal{C},E_{\mathcal{R}})$triangleright Theorem B.9
@@
Proof 5.1.
The $(2,\mathcal{O}(\log\Delta)\cap\mathcal{O}(\log\log n))$-ruling set can be computed in $\mathcal{O}(\log\log n)$ rounds according to Corollary B.5. Hence, each cluster has diameter at most $\mathcal{O}(\log\log n)$ and can be computed in $\mathcal{O}(\log\log n)$ rounds as well. For the $2$-edge ruling set the deterministic algorithm from Theorem B.7 running in $\mathcal{O}(\log^{\ast}n)$ rounds is still fast enough for our purposes here. According to Theorem B.9 and Section 4.3 the hypergraph sinkless orientation can be computed in $\mathcal{O}(\log_{\Delta}\log n)$ rounds on the cluster graph, where we incur a $\mathcal{O}(\log\Delta)$-round simulation overhead. Finally, the extension to the edges inside each cluster can be computed in $\mathcal{O}(\log\log n)$ rounds. Hence, the total runtime works out to be
[…]
Similar to the deterministic case we could also get a $\mathcal{O}(\log_{\Delta}\log n)$-round reduction to randomized reduction via replacing $(2,\mathcal{O}(\log\Delta)\cap\mathcal{O}(\log\log n))$-ruling set by an MIS. For randomized algorithms there is however a gap between the complexities for MIS and $(2\Delta-1)$-edge coloring. While $(2\Delta-1)$-edge coloring can be solved in $\widetilde{\mathcal{O}}(\log^{5/3}\log n)$ rounds [20], applying the shattering framework of [18] to the deterministic MIS algorithm from [20] takes $\mathcal{O}(\log\Delta)+\widetilde{\mathcal{O}}(\log^{8/3}\log n)$ rounds. Hence, this alternative reduction does not yield any useful improvements at the moment.
6 Conclusion
Here we discuss some open questions and limits of our approach.
Why can we not use this approach for $\Delta$-vertex coloring?
The existing deterministic $\Delta$-coloring algorithms from the literature [37, 38, 22, 25] all follow a similar three-step approach: First they compute a ruling set $R$, then they color nodes outside $R$ in order of decreasing distance from $R$ and finally they color each vertex $v\in R$ via recoloring only a small-diameter subgraph around $v$. If we try to translate our approach into the vertex coloring setting we see that it is still possible to formulate an analogous condition for the extendability of partial vertex colorings to star graphs. However, clusters of larger diameter cannot be dealt with a similar approach. An intuitive reason for this is that the $\Delta$-vertex coloring is a much more restrictive problem. In order to find a color for the root $r$, it needs to have two neighbors $v_{1},v_{2}$ with the same color. In the literature this is often called having “slack” at $r$. If we now consider a $\Delta$-regular tree we unfortunately see that this conditions propagates through the layers of the tree, since it completely determines the colors of all vertices in the subtrees rooted at $v_{1}$ or $v_{2}$, up to permutations. One reason this approach works better for edge-coloring is the fact that one has just a bit more freedom in how one can place slack: If we want to color a tree from the leaves to the root, in the vertex coloring case, we need to place two nodes with the same color in layer $k+1$ to get a node with slack in layer $k$. In contrast, to obtain an edge with slack in layer $k$ we can use one edge from layer $k+1$ and the second one from layer $k$.
Appendix A Deferred proofs
A.1 Replacing the maximal matching by a $2$-edge ruling set
In this section we show that we can replace the maximal matching $M$ by a cheaper $2$-edge ruling set $E_{\mathcal{R}}$ of $G-E_{T}$. In this case, every leaf-vertex of a cluster sees at least one ruling set edge in its $3$-hop neighborhood. Here we need to be a little bit more careful in how we set up our hypergraph. Firstly, each vertex may only send one proposal to one of its closest ruling set edges. Secondly, each vertex may only propose to a ruling set edge $e\in E_{\mathcal{R}}$, if all the vertices on a shortest path towards this have proposed to $e$ as well. For an illustration we refer to Figure 4.
Figure 4: The ruling set edges are marked orange and dashed lines indicate the proposals.
Definition A.1 (Proposal scheme).
Definition A.2.
  • If $v$ is incident to $e$, then we simply add $e$ to $M^{\prime}$.
Now we prove an analogous result to Lemma 4.14.
Lemma A.3.

### v2

@@ 4.3 Phase 2: Assign two exclusive edges to each cluster
In this phase, we rearrange the edges in the maximal matching $M$ in order to assign two exclusive edges to each cluster, while preserving the matching property. These exclusive edges are essential for adjusting the coloring of the intercluster edges, as laid out in Lemma 4.16. This step relies on the HSO subroutine from Theorem A.8, and we now detail the specific instance used in our algorithm. The high level idea is that each cluster sends proposals to nearby edges in the matching $M$. Each edge accepts exactly one proposal and we swap the edge towards the cluster that won it. This process can be modeled as an HSO instance where each matching edge in $M$ represents an hyperedge that contains as vertices all clusters that sent a proposal to that edge. Since a directed hyperedge $e$ is outgoing for exactly one of its vertices $v\in e$, we will say that $v$ won the proposal for $e$.
However, the hypergraph sinkless orientation problem only guarantees at least one outgoing hyperedge for each vertex. Therefore, in order to achieve two exclusive hyperedges for each cluster, we simply split each cluster into two virtual nodes and let both participate in the HSO independently. In the following we formalize this idea and prove that this provides the desired properties as well as a deterministic runtime of $\mathcal{O}(\log n)$ rounds. First, we let each cluster $C^{-1}(r)\in\mathcal{C}$ that is not already edge-degree choosable select a subset of vertices $S\subseteq V(C^{-1}(r))$ to send proposals to a matching edges in their neighborhood. Let $\mathcal{C}^{\prime}$ denote the set of clusters, which are not already edge-degree choosable. Crucially, in order to satisfy the colorful condition from Lemma 3.4 we require that the vertices in $S$ are all leaves in the spanning tree $T(r)\subseteq C^{-1}(r)$ at the same distance from the root $r$.
@@
For each cluster $C^{-1}(r)\in\mathcal{C}^{\prime}$, there exists an integer $k>0$ and a subset $S\subseteq V(C^{-1}(r))$ of size at least $4\cdot\Delta^{3}$ such that every vertex $v\in S$ is a leaf in the tree $T(r)$ that satisfies $\mathrm{dist}(r,v)=k$.
@@
To compute an orientation with two outgoing edges per cluster we simply split each vertex into two virtual nodes with half the degree each. Thus, the minimum degree of the resulting hypergraph $H^{\prime}$ is still at least $\Delta^{2}$. Therefore, we can solve HSO in $\mathcal{O}(\log_{\delta/r}n)=\mathcal{O}(\log_{\Delta}n)$ rounds deterministically and in $\mathcal{O}(\log_{\delta/r}\delta+\log_{\delta/r}\log n)=\mathcal{O}(\log_{\Delta}\log n)$ rounds randomized using the algorithms from Theorem A.8 and Theorem A.9.
@@ 4.4 Phase 3: Switch colors in order to complete the coloring
Now that each cluster has been assigned two exclusive edges located at the same distance from its central vertex, we show how to modify their colors to ensure that the colorful condition from Lemma 3.4 is satisfied for every cluster. To prove that this procedure can be applied to all clusters in parallel, we argue that the colorful condition can still be fulfilled even if the colors of all edges not adjacent to the two exclusive edges are modified adversarially.
@@
Recall that $V_{k}:=\{v\in V_{T}:\mathrm{dist}_{T}(v,r)=k\}$. Since $u$ and $u^{\prime}$ are leaves in $T$, they are incident to exactly $\Delta-1$ edges in $E\setminus E_{T}$. Hence, $\varphi(N_{E}(V_{k}))$ already contains at least $\Delta-1$ distinct colors. If $\lvert\varphi(N_{E}(V_{k}))\rvert\geq\Delta$ we simply set $\varphi^{\prime}=\varphi$. Otherwise, if $\lvert\varphi(N_{E}(V_{k}))\rvert=\Delta-1$, we set $\varphi^{\prime}(uv)=2\Delta-2$. Since $\varphi(e)<2\Delta-2$ for all $e\in E_{k}$ we only need to argue that the color $\varphi(uv)$ still remains in $\varphi^{\prime}(E_{k})$. Consider the sets $N:=\{uw:uw\in E\setminus E_{T}\}$ and $N^{\prime}:=\{u^{\prime}w:u^{\prime}w\in E\setminus E_{T}\}$. Since $u,u^{\prime}$ are both leaves in $T$, both $N$ and $N^{\prime}$ have cardinality $\Delta-1$. Moreover, since $\lvert\varphi(E_{k})\rvert=\Delta-1$ and any two edges in $N$ (or $N^{\prime}$) are adjacent, we have that $\varphi(N)=\varphi(N^{\prime})$. Hence, there exists an edge $u^{\prime}v^{\prime\prime}$ such that $\varphi(uv)=\varphi(u^{\prime}v^{\prime\prime})=\varphi^{\prime}(u^{\prime}v^{\prime\prime})\in\varphi^{\prime}(E_{k})$. Thus, we have $\varphi^{\prime}(E_{k})=\Delta$. Note that this argument holds even if the colors of the edges not adjacent to $u$ or $v$ are modified adversarially.
4.5 Proof of Theorem 1.2
See 1.2
Proof 4.18.
@@
For $\Delta=\mathcal{O}(1)$ we apply the $(3\Delta/2)$-edge coloring algorithm from Theorem A.13 to get a $(2\Delta-2)$-edge coloring in $\mathcal{O}(\log n)$ rounds. Hence, for the remainder of the proof we assume $\Delta\geq\Delta_{0}$ for a sufficiently large constant $\Delta_{0}$. Since the power graph $G^{8}$ has $n$ vertices and maximum degree polynomial in $\Delta$, an MIS can be computed on it in time $T_{\mathrm{MIS}}(n,\mathrm{poly}(\Delta))$. Next, the clustering is possible in a constant number of rounds, since the radius of each cluster is at most 8. Also, both computing the $(2\Delta-3)$-edge coloring $\varphi$ of $G-E_{T}$ and finding a maximal matching in $E_{T}$ can be reduced to MIS, and thus can be solved in time $T_{\mathrm{MIS}}(\Delta^{2}\cdot n,\mathrm{poly}(\Delta))$. Then, according to Section 4.3 we can modify the matching in a way that each cluster receives two exclusive edges. The rearrangement can be modeled as an instance of HSO with minimum degree $\frac{\Delta^{3}}{4}$ and maximum rank $\Delta^{2}$. Since each cluster has constant diameter, simulating this hypergraph incurs just a constant overhead factor. Thus, the deterministic HSO algorithm from Theorem A.8 runs in $\mathcal{O}(\log_{\Delta}n)$ rounds. Next, we apply Lemma 4.16 to ensure that for each cluster there is a $k\leq 8$ such that $\lvert\varphi(N_{E}(V_{k}))\rvert\geq\Delta$. By Lemma 3.4 we can extend $\varphi$ to a proper $(2\Delta-2)$-edge coloring of $G$.
4.6 Replacing MIS and maximal matching (Proof of Theorem 1.1)
@@
In this subsection we show that computing an MIS is not strictly necessary for our reduction. Indeed, using a $(2,k)$-ruling set changes only the diameter of our clusters. Since increasing the cluster diameter increases the overhead for simulating the auxiliary hypergraph, this induces a tradeoff between $k$ and the runtime of computing a $(2,k)$-ruling set. Setting $k=\mathcal{O}(\log\Delta)$ turns out to be the best choice for our purposes. Secondly, we show how to substitute the maximal matching by a $2$-edge ruling set. This change is technically more involved, as it requires us to generalize the definition of our auxiliary hypergraph. The necessary changes to Algorithm 1 are also highlighted red in Algorithm 3.
@@
Algorithm 3 Reduction from $(2\Delta-2)$-edge coloring to $(2\Delta-1)$-edge coloring (high level overview)
1:Compute a $(2,\mathcal{O}(\log\Delta))$-ruling set $\mathcal{R}$ on $G^{8}$triangleright Corollary A.3
@@
4:Compute a $2$-edge ruling set $E_{\mathcal{R}}$ of $T$triangleright Theorem A.7
5:Compute $2$-HSO $\psi:\mathcal{C}\to E_{\mathcal{R}}$ in $H=(\mathcal{C},E_{\mathcal{R}})$triangleright Theorem A.8
@@
10:Extend $\varphi$ to a $(2\Delta-2)$-edge coloring $\varphi^{\star}$ of $G$triangleright Lemma 3.4
Secondly, we show that we can replace the maximal matching $M$ by a cheaper $2$-edge ruling set $E_{\mathcal{R}}$ of $G-E_{T}$. In this case, every leaf-vertex of a cluster sees at least one ruling set edge in its $3$-hop neighborhood. Here we need to be a little bit more careful in how we set up our hypergraph. Firstly, each vertex may only send one proposal to one of its closest ruling set edges. Secondly, each vertex may only propose to a ruling set edge $e\in E_{\mathcal{R}}$, if all the vertices on a shortest path towards this have proposed to $e$ as well. For an illustration we refer to Figure 6.
(a) Each vertex proposes to one of the nearest matching edges (orange) within its two-hop neighborhood. Note that vertices lying on the shortest path to a matching edge will propose to the same edge. These proposals are represented by dashed arrows. According to the HSO result, each orange matching edge accepts exactly one proposal, which is indicated by a blue arrow.
(b) Based on the HSO result (blue arrows), the matching is shifted toward the vertices that won the matching edge. The first edge along the shortest path to these vertices (orange) becomes the new matching edge.
Figure 6: Illustration of a simplified example showing how the HSO is modeled as a proposal instance for assigning clusters two exclusive edges.
Definition 4.19 (Proposal scheme).
Definition 4.20.
  • If $v$ is incident to $e$, then we simply add $e$ to $M^{\prime}$.
Now we prove an analogous result to Lemma 4.14.
Lemma 4.21.
Proof 4.22.
See 1.1
Proof 4.23.
The $(2,\mathcal{O}(\log\Delta))$-ruling set can be computed in $\mathcal{O}(\log\Delta+\log^{\ast}n)$ rounds according to Corollary A.3. As an immediate consequence, the diameter of our clusters now increases to $\mathcal{O}(\log\Delta)$. This implies that simulating the hypergraph now incurs a $\mathcal{O}(\log\Delta)$ factor overhead. Therefore the complexity of computing a HSO increases to $\mathcal{O}(\log_{\Delta}n\cdot\log\Delta)=\mathcal{O}(\log n)$ rounds. The $(2\Delta-3)$-edge coloring of $G-E_{T}$ remains unchanged. Further, we can also relax the requirements for our matching $M$ to $2$-edge ruling set, which can be computed in $\mathcal{O}(\log^{\ast}n)$ rounds. Therefore the total runtime of our reduction works out to be
@@
Proof 4.24.
For the first claim we plug in the $\mathcal{O}(\log^{12}\Delta+\log^{\ast}n)$-round algorithm from Theorem A.12 into Theorem 1.1. For the latter we plug in the $\widetilde{\mathcal{O}}(\log^{5/3}n)$-round MIS algorithm from Theorem A.10 into Theorem 1.2 instead.
@@ 5 Randomized $(2\Delta-2)$-edge coloring algorithm
In order to achieve a $\mathcal{O}(\log\log n)$-round randomized reduction, we only need to replace the ruling set algorithm by a faster randomized variant in Corollary A.5 and the deterministic HSO algorithm by its randomized counterpart in Theorem A.9. The changes to the optimized deterministic version are also highlighted red in Algorithm 4.
@@
1:Compute a $(2,\mathcal{O}(\log\Delta)\cap\mathcal{O}(\log\log n))$-ruling set $\mathcal{R}$ on $G^{8}$triangleright Corollary A.5
@@
4:Compute a $2$-edge ruling set $E_{\mathcal{R}}$ of $T$triangleright Theorem A.7
5:Compute $2$-HSO $\psi:\mathcal{C}\to E_{\mathcal{R}}$ in $H=(\mathcal{C},E_{\mathcal{R}})$triangleright Theorem A.9
@@
See 1.4
Proof 5.1.
The $(2,\mathcal{O}(\log\Delta)\cap\mathcal{O}(\log\log n))$-ruling set can be computed in $\mathcal{O}(\log\log n)$ rounds according to Corollary A.5. Hence, each cluster has diameter at most $\mathcal{O}(\log\log n)$ and can be computed in $\mathcal{O}(\log\log n)$ rounds as well. For the $2$-edge ruling set the deterministic algorithm from Theorem A.7 running in $\mathcal{O}(\log^{\ast}n)$ rounds is still fast enough for our purposes here. According to Theorem A.9 and Section 4.3 the hypergraph sinkless orientation can be computed in $\mathcal{O}(\log_{\Delta}\log n)$ rounds on the cluster graph, where we incur a $\mathcal{O}(\log\Delta)$-round simulation overhead. Finally, the extension to the edges inside each cluster can be computed in $\mathcal{O}(\log\log n)$ rounds. Hence, the total runtime works out to be
[…]
@@
Proof 5.2.
To prove the first claim we plug in the deterministic $\mathcal{O}(\log^{12}\Delta+\log^{\ast}n)$-round algorithm from Theorem A.12 into Theorem 1.4, while the second claim uses the randomized $\widetilde{\mathcal{O}}(\log^{5/3}\log n)$-round $(2\Delta-1)$-edge coloring algorithm from Corollary A.11.
Similar to the deterministic case we could also get a $\mathcal{O}(\log_{\Delta}\log n)$-round reduction to MIS via replacing the $(2,\mathcal{O}(\log\Delta)\cap\mathcal{O}(\log\log n))$-ruling set by an MIS. However, applying the shattering framework of [35] to the deterministic MIS algorithm from [37] still takes $\mathcal{O}(\log\Delta)+\widetilde{\mathcal{O}}(\log^{8/3}\log n)$ rounds. Hence, the additional cost of computing an MIS currently outweighs the time saved from the faster reduction.
Appendix A Complementary subroutines
A.1 Ruling sets
Theorem A.1 (ruling set algorithm [34, Theorem 8]).
Theorem A.2 (ruling set algorithm [67, Theorem 3]).
Corollary A.3.